# Balanced rectangles over Sturmian words and minimal discrepancy intervals

### v1

@@ 5.4.  Minimising distances in certain ranges of integers
(5.14) $k_{\geq M}(n^{\prime})\equiv M+1\pmod{2}$.
@@
We are left with the case $(q_{T}-n)^{[\leq M-1]}=0$. Then $k_{0}(q_{T}-n)\geq M+1$. Lemma 5.1(2) implies $\lVert(q_{T}-n)\alpha\rVert<\lVert q_{M}\alpha\rVert$, and so $\lVert n\alpha\rVert<\lVert q_{M}\alpha\rVert+\lVert q_{T}\alpha\rVert\leq\lVert q_{M}\alpha\rVert+\lVert q_{N+2}\alpha\rVert$. But then Lemma 5.1(3) implies that $k_{0}(n)\geq M$, which contradicts the assumption $n^{[\leq M-1]}>0$. ∎
Finally, we combine the previous two lemmas in the way we will want to apply them later.
@@
Let $q_{M}<n<q_{N}$ for some $M\geq 2$ (or $M\geq 1$ if $\alpha<1/2$). Assume that $b_{M}(n)>0$, and that $n$ is not of the shape $n=q_{M}+\sum_{k=M+1+2t}^{N}b_{k}q_{k}$. Then for $T\geq N+2$ we have $k_{\geq M}(q_{T}-n)\equiv M\pmod{2}$.
@@
If $n^{[M-1]}>0$, then Lemma 5.17 implies that $k_{\geq M}(q_{T}-n)\equiv k_{\geq M}(n)=M\pmod{2}$. If $n^{[M-1]}=0$, then $k_{0}(n)=M$, and so Lemma 5.16 implies that $k_{\geq M}(q_{T}-n)\equiv M\pmod{2}$. ∎
@@ 6.  Proof of the full characterisation
In this section, we prove our main result, namely the full characterisation of balanced rectangles of Sturmian sequences (Theorem 2.1). As usual, let $\alpha\in(0,1)$ be irrational and fix integers $2\leq m\leq n$. In view of Theorem 3.1, we can phrase Theorem 2.1 in terms of balanced intervals. Moreover, as mentioned in Remark 4.5, we can use Lemma 4.3, setting
$A$ $:=\{\xi_{0},\xi_{1},\ldots,\xi_{m-1}\}\quad\text{with}\quad\xi_{\ell}:=\left\{\ell\alpha\right\}\text{ for }0\leq\ell\leq m-1;$
@@
This means that we need to characterise when $f_{\textnormal{left}}$ and $f_{\textnormal{right}}$ from Definition 4.2 are bijective (and we know from Lemma 4.3 that one is bijective if and only if the other is bijective, so we can focus on either of them).
Recall that to describe $f_{\textnormal{left}},f_{\textnormal{right}}$, we need to find the closest points from the set $A$ on either side of $\xi_{\ell}+\delta\bmod 1=\left\{(n+\ell)\alpha\right\}$. In other words, we need to subtract some positive integer $x$ from $n+\ell$, so that $n+\ell-x$ falls into [0,m1][0,m-1], and so that $\left\{x\alpha\right\}$ causes a minimal shift to the left or to the right.
@@
(6.2) $x_{\textnormal{left}}(\ell),x_{\textnormal{right}}(\ell)\in[n+\ell-m+1,n+\ell]=:X(\ell).$
@@
and we can check the formula for $x_{\textnormal{right}}(\ell)$ analogously. ∎
@@
Let $(f,x)=(f_{\textnormal{left}},x_{\textnormal{left}})$ or $(f_{\textnormal{right}},x_{\textnormal{right}})$. Then $f$ is bijective if and only if
$x(\ell)-\ell\neq x(\ell^{\prime})$-$\ell^{\prime}$
@@
This follows directly from the definition: We have $x(\ell)=n+\ell-f(\ell)$, which gives us $f(\ell)=n+\ell-x(\ell)$, and so $f(\ell)=f(\ell^{\prime})\iff x(\ell)-\ell=x(\ell^{\prime})$-$\ell^{\prime}$. Since $f$ is a map on the finite set $\{0,1,\ldots,m-1\}$, it is injective if and only if it is bijective. Thus, $f$ is bijective if and only if $f(\ell)\neq f(\ell^{\prime})$ for all $\ell\neq\ell^{\prime}$, and by the previous argument this is equivalent to $x(\ell)-\ell\neq x(\ell^{\prime})$-$\ell^{\prime}$ for all $\ell\neq\ell^{\prime}$. ∎
A brief recap: We want to characterise all $2\leq m\leq n$ such that the $m\times n$ rectangles of the Sturmian words with slope $\alpha$ are balanced (prove Theorem 2.1). The $m\times n$ rectangles of the Sturmian words with slope $\alpha$ being balanced is equivalent to the intervals of length $\left\{n\alpha\right\}$ being balanced with respect to $(\alpha,m)$ (Theorem 3.1). This again is equivalent to the function $f=f_{\textnormal{left}}$ or $f=f_{\textnormal{right}}$ being bijective (Lemma 4.3). And to decide whether $f$ is bijective, we can use Lemma 6.3 and the corresponding function $x=x_{\textnormal{left}}$ or $x=x_{\textnormal{right}}$, which can be defined via Lemma 6.1. Indeed, this is our strategy. In order to unify our arguments, recall that in the previous section we defined
@@
In the rest of this section, we will always either set $(f,x,\langle\cdot\rangle)=(f_{\textnormal{left}},x_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{left}})$ or $(f,x,\langle\cdot\rangle)=(f_{\textnormal{right}},x_{\textnormal{right}},\langle\cdot\rangle_{\textnormal{right}})$. In either case, we can now phrase Lemma 6.1 as
$\langle x(\ell)\alpha\rangle=\min_{x\in X(\ell)}\langle x\alpha\rangle,$
where the range $X(\ell)=[n+\ell-m+1,\leq n+\ell]$ was defined in (6.2). In view of this and Lemma 6.3, we are interested in integers $x$ in the range
@@
Let $x_{\textnormal{left}}^{*},x_{\textnormal{right}}^{*}$ be the integers that minimise $\langle x\alpha\rangle_{\textnormal{left}}$ or $\langle x\alpha\rangle_{\textnormal{right}}$, respectively, in the full range [nm+1,n+m1][n-m+1,n+m-1]. Again, according to the setting, we will write just $x^{*}$ for $x_{\textnormal{left}}^{*}$ or $x_{\textnormal{right}}^{*}$. In other words, in our notation we have
$\langle x^{*}\alpha\rangle=\min_{n-m+1\leq x\leq n+m-1}\langle x\alpha\rangle.$
Note that the only integer that occurs in the range $X(\ell)=[n+\ell-m+1,n+\ell]$ for every $0\leq\ell\leq m-1$ is the integer $n$. Therefore, the case $x^{*}=n$ is particularly easy, and we start with this case.
6.1.  The case $n=x^{*}$
@@
Let $(f,x^{*})=(f_{\textnormal{left}},x_{\textnormal{left}}^{*})$ or $(f,x^{*})=(f_{\textnormal{right}},x_{\textnormal{right}}^{*})$. If $x^{*}=n$, then $f$ is bijective.
@@
Since $x^{*}=n\in X(\ell)=[n+\ell-m+1,n+\ell]$ for every $0\leq\ell\leq m-1$, we have in fact $x(\ell)=x^{*}$ for all $0\leq\ell\leq m-1$. Thus $x(\ell)-\ell=x^{*}-\ell\neq x^{*}-\ell^{\prime}=x(\ell^{\prime})$-$\ell^{\prime}$ for all $\ell\neq\ell^{\prime}$, and by Lemma 6.3 the function $f$ is bijective. ∎
We also characterise the case $x^{*}=n$ in terms of the Ostrowski representations of $m,n$.
@@
Let $2\leq m\leq n$ with $q_{M-1}<m\leq q_{M}$. We have $n\in\{x_{\textnormal{left}}^{*},x_{\textnormal{right}}^{*}\}$ if and only if the Ostrowski representation of $n$ with respect to $\alpha$ is of one of the following two shapes:
@@
If $\langle n\alpha\rangle>1/2$ for $\langle\cdot\rangle=\langle\cdot\rangle_{\textnormal{left}}$ or $\langle\cdot\rangle=\langle\cdot\rangle_{\textnormal{right}}$, then, by Lemma 5.4, one of $\langle(n-1)\alpha\rangle,\langle(n+1)\alpha\rangle$ must be strictly smaller than $\langle n\alpha\rangle$. Since n1,n+1n-1,n+1 are both in the range [nm+1,n+m1][n-m+1,n+m-1], we have $n\neq x^{*}$. Therefore, it suffices to consider the choice of $\langle\cdot\rangle\in\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$ for which $\langle n\alpha\rangle<1/2$, and check whether $x^{*}=n$. We systematically go through all possible representations of $n$.
Case 1: Either $L\geq M$, or L=M1L=M-1 and $n=q_{M-1}+\sum_{k=M+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{M+2t}\neq 0$. In other words, $n$ has the representation from (a) or (b). Then Lemma 5.12 says that $n$ minimises $\langle n\alpha\rangle$ in the range $n-q_{M}+1,n+q_{M}-1]\supseteq[n-m+1,n+m-1]$. Thus, we have indeed $n=x^{*}$.
Case 2: L=M1L=M-1 and $n=q_{M-1}+\sum_{k=M+1+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{M+1+2t}\neq 0$. Set $n_{1}:=\sum_{k=M+1+2t}^{N}b_{k}q_{k}$. By Lemma 5.2 we have $\langle n_{1}\alpha\rangle<1/2$ as well, except possibly if M1=0M-1=0 or M1=1M-1=1. In the exceptional cases (i.e., if $\left\{n\alpha\right\}$ is “on the wrong side of 1/21/2”), we can apply Lemma 5.3 and obtain $\langle(n+1)\alpha\rangle<\langle n\alpha\rangle$. Otherwise, if $\langle n_{1}\alpha\rangle<1/2$, then since $q_{M-1}<m$, the integer $n_{1}$ is in the range [nm+1,n+m1][n-m+1,n+m-1]. By Lemma 5.6, we have $\langle n_{1}\alpha\rangle<\langle n\alpha\rangle$, and so $n\neq x^{*}$.
Case 3: L=M1L=M-1 and $n=b_{M-1}q_{M-1}+\sum_{k=M}^{N}b_{k}q_{k}$ with $b_{M-1}\geq 2$. If $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$), we set $n_{1}:=n-q_{M-1}>n-m$. Then, by Lemma 5.7, we have $\langle n_{1}\alpha\rangle<\langle n\alpha\rangle$, and so $n\neq x^{*}$. If $L=1$ (or $L=0$ if $\alpha<1/2$), the same follows from Lemma 5.11.
Case 4: $L\leq M-2$. If $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$), we set $n_{1}:=n+q_{L+1}<n+m$. Then by Lemma 5.9, we have $\langle n_{1}\alpha\rangle<\langle n\alpha\rangle$, and so $n\neq x^{*}$. If $L=1$ (or $L=0$ if $\alpha<1/2$), the same follows from Lemma 5.10. ∎
The two above lemmas give us a partial result towards the full characterisation in Theorem 2.1; note that the cases (a) and (b) in Lemma 6.5 are almost the same as the cases (i) and (ii) in Theorem 2.1. Next, we provide some lemmas that will be useful in the situation where $x^{*}\neq n$. We start with two lemmas which will, roughly speaking, allow us to assume $x^{*}<n$ without loss of generality.
6.2.  Switching beween $n$ and $q_{T}-n$

### v2

@@ 5.4.  Minimising distances in certain ranges of integers
(5.14) $k_{\geq M}(n^{\prime\prime})\equiv M+1\pmod{2}$.
@@
We are left with the case $(q_{T}-n)^{[\leq M-1]}=0$. Then $k_{0}(q_{T}-n)\geq M+1$. Lemma 5.1(2) implies $\lVert(q_{T}-n)\alpha\rVert<\lVert q_{M}\alpha\rVert$, and so $\lVert n\alpha\rVert<\lVert q_{M}\alpha\rVert+\lVert q_{T}\alpha\rVert\leq\lVert q_{M}\alpha\rVert+\lVert q_{N+2}\alpha\rVert$. But then Lemma 5.1(3) implies that $k_{0}(n)\geq M$, which contradicts the assumption in the lemma that $n^{[\leq M-1]}>0$. ∎
We combine the previous two lemmas in the way we will want to apply in the next section.
@@
Let $\alpha\in(0,1/2)$ and $q_{M}<n<q_{N-1}$ for some $M\geq 1$. Assume that $b_{M}(n)>0$, and that $n$ is not of the shape $n=q_{M}+\sum_{k=M+1+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{M+1+2t}>0$. Then for $T\geq N+2$ we have $k_{\geq M}(q_{T}-n)\equiv M\pmod{2}$.
@@
If $n^{[\leq M-1]}>0$, then Lemma 5.16 implies that $k_{\geq M}(q_{T}-n)\equiv k_{\geq M}(n)=M\pmod{2}$. If $n^{[\leq M-1]}=0$, then $k_{0}(n)=M$, and so Lemma 5.15 implies that $k_{\geq M}(q_{T}-n)\equiv M\pmod{2}$. ∎
@@ 6.  Proof of the full characterisation
In this section, we prove our main result, namely the full characterisation of balanced rectangles of Sturmian sequences (Theorem 2.1). As before, let $\alpha\in(0,1/2)$ be irrational and fix integers $2\leq m\leq n$. In view of Theorem 3.1, we can phrase Theorem 2.1 in terms of balanced intervals. Moreover, as mentioned in Remark 4.5, we can use Lemma 4.3, setting
$B$ $:=\{\xi_{0},\xi_{1},\ldots,\xi_{m-1}\}\quad\text{with}\quad\xi_{\ell}:=\left\{\ell\alpha\right\}\text{ for }0\leq\ell\leq m-1;$
@@
This means that we need to characterise the situations when $f_{\textnormal{left}}$ and $f_{\textnormal{right}}$ from Definition 4.2 are bijective (and we know from Lemma 4.3 that one is bijective if and only if the other is bijective, so we can focus on either of them).
Recall that to describe $f_{\textnormal{left}},f_{\textnormal{right}}$, we need to find the closest points from the set $B$ on either side of $\xi_{\ell}+\delta\bmod 1=\left\{(n+\ell)\alpha\right\}$. In other words, we need to subtract some positive integer $x$ from $n+\ell$, so that $n+\ell-x$ falls into [0,m1][0,m-1], and so that $\left\{x\alpha\right\}$ causes a minimal shift to the left or to the right.
@@
(6.2) $x_{\textnormal{left}}(\ell),x_{\textnormal{right}}(\ell)\in[n+\ell-m+1,n+\ell]=:X(\ell)$
for $0\leq\ell\leq m-1$.
@@
and we can check the formula for $\left\{x_{\textnormal{right}}(\ell)\alpha\right\}$ analogously. ∎
@@
Let $(f_{*},x_{*})=(f_{\textnormal{left}},x_{\textnormal{left}})$ or $(f_{\textnormal{right}},x_{\textnormal{right}})$. Then $f_{*}$ is bijective if and only if
$x_{*}(\ell)-\ell\neq x_{*}(\ell^{\prime})$-$\ell^{\prime}$
@@
This follows directly from the definition: We have $x_{*}(\ell)=n+\ell-f_{*}(\ell)$, which gives us $f_{*}(\ell)=n+\ell-x_{*}(\ell)$, and so $f_{*}(\ell)=f_{*}(\ell^{\prime})\iff x_{*}(\ell)-\ell=x_{*}(\ell^{\prime})$-$\ell^{\prime}$. Since $f_{*}$ is a map on the finite set $\{0,1,\ldots,m-1\}$, it is injective if and only if it is bijective. Thus, $f_{*}$ is bijective if and only if $f_{*}(\ell)\neq f_{*}(\ell^{\prime})$ for all $\ell\neq\ell^{\prime}$, and by the previous argument this is equivalent to $x_{*}(\ell)-\ell\neq x_{*}(\ell^{\prime})$-$\ell^{\prime}$ for all $\ell\neq\ell^{\prime}$. ∎
A brief recap: We want to characterise all $2\leq m\leq n$ such that the $m\times n$ rectangles of the Sturmian words with slope $\alpha$ are balanced (prove Theorem 2.1). The $m\times n$ rectangles of the Sturmian words with slope $\alpha$ being balanced is equivalent to the intervals of length $\left\{n\alpha\right\}$ being balanced with respect to $(\alpha,m)$ (Theorem 3.1). This again is equivalent to the function $f_{*}=f_{\textnormal{left}}$ or $f_{*}=f_{\textnormal{right}}$ being bijective (Lemma 4.3). And to decide whether $f_{*}$ is bijective, we can use Lemma 6.3 and the corresponding function $x_{*}=x_{\textnormal{left}}$ or $x_{*}=x_{\textnormal{right}}$, which can be defined via Lemma 6.1. Indeed, this is our strategy. In order to unify our arguments, recall that in the previous section we defined
@@
In the rest of this section, we will always either set $(f_{*},x_{*},\langle\cdot\rangle_{*})=(f_{\textnormal{left}},x_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{left}})$ or $(f_{*},x_{*},\langle\cdot\rangle_{*})=(f_{\textnormal{right}},x_{\textnormal{right}},\langle\cdot\rangle_{\textnormal{right}})$. In either case, we can now phrase Lemma 6.1 as
$\langle x_{*}(\ell)\alpha\rangle_{*}=\min_{x\in X(\ell)}\langle x\alpha\rangle_{*},$
where the range $X(\ell)=[n+\ell-m+1,n+\ell]$ was defined in (6.2). In view of this and Lemma 6.3, we are interested in integers $x$ in the range
@@
Let $\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}$ be the integers that minimise $\langle x\alpha\rangle_{\textnormal{left}}$ or $\langle x\alpha\rangle_{\textnormal{right}}$, respectively, in the full range [nm+1,n+m1][n-m+1,n+m-1]. Again, according to the setting, we will write just $\widehat{x}_{*}$ for $\widehat{x}_{\textnormal{left}}$ or $\widehat{x}_{\textnormal{right}}$. In other words, in our notation we have
$\langle\widehat{x}_{*}\alpha\rangle_{*}=\min_{n-m+1\leq x\leq n+m-1}\langle x\alpha\rangle_{*}.$
Note that the only integer that occurs in the range $X(\ell)=[n+\ell-m+1,n+\ell]$ for every $0\leq\ell\leq m-1$ is the integer $n$. Therefore, the case $\widehat{x}_{*}=n$ is particularly easy, and we start with this case.
6.1.  The case $n=\widehat{x}_{*}$
@@
Let $(f_{*},\widehat{x}_{*})=(f_{\textnormal{left}},\widehat{x}_{\textnormal{left}})$ or $(f_{*},\widehat{x}_{*})=(f_{\textnormal{right}},\widehat{x}_{\textnormal{right}})$. If $\widehat{x}_{*}=n$, then $f_{*}$ is bijective.
@@
Since $\widehat{x}_{*}=n\in X(\ell)=[n+\ell-m+1,n+\ell]$ for every $0\leq\ell\leq m-1$, we have in fact $x_{*}(\ell)=\widehat{x}_{*}$ for all $0\leq\ell\leq m-1$. Thus, $x_{*}(\ell)-\ell=\widehat{x}_{*}-\ell\neq\widehat{x}_{*}-\ell^{\prime}=x_{*}(\ell^{\prime})$-$\ell^{\prime}$ for all $\ell\neq\ell^{\prime}$, and by Lemma 6.3 the function $f_{*}$ is bijective. ∎
We also characterise the case $\widehat{x}_{*}=n$ in terms of the Ostrowski representations of $m,n$.
@@
Let $2\leq m\leq n$ with $q_{M-1}<m\leq q_{M}$. We have $n\in\{\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}\}$ if and only if the Ostrowski representation of $n$ with respect to $\alpha$ is of one of the following two shapes:
@@
If $\langle n\alpha\rangle_{*}>1/2$ for $\langle\cdot\rangle_{*}=\langle\cdot\rangle_{\textnormal{left}}$ or $\langle\cdot\rangle_{*}=\langle\cdot\rangle_{\textnormal{right}}$, then, by Lemma 5.4, one of $\langle(n-1)\alpha\rangle_{*},\langle(n+1)\alpha\rangle_{*}$ must be strictly smaller than $\langle n\alpha\rangle_{*}$. Since n1,n+1n-1,n+1 are both in the range [nm+1,n+m1][n-m+1,n+m-1], we have $n\neq\widehat{x}_{*}$. Therefore, it suffices to consider the choice of $\langle\cdot\rangle_{*}\in\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$ for which $\langle n\alpha\rangle_{*}<1/2$, and check whether $\widehat{x}_{*}=n$. We systematically go through all possible representations of $n$.
Case 1: Either $L\geq M$, or L=M1L=M-1 and $n=q_{M-1}+\sum_{k=M+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{M+2t}\neq 0$. In other words, $n$ has the representation from (a) or (b). Then Lemma 5.12 says that $n$ minimises $\langle n\alpha\rangle_{*}$ in the range $n-q_{M}+1,n+q_{M}-1]\supseteq[n-m+1,n+m-1]$. Thus, we have indeed $n=\widehat{x}_{*}$.
Case 2: L=M1L=M-1 and $n=q_{M-1}+\sum_{k=M+1+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{M+1+2t}\neq 0$. Set $n_{1}:=\sum_{k=M+1+2t}^{N}b_{k}q_{k}$. By Lemma 5.2 we have $\langle n_{1}\alpha\rangle_{*}<1/2$ as well, except possibly if M1=0M-1=0. In the exceptional case (i.e., if $\left\{n\alpha\right\}$ is “on the wrong side of 1/21/2”), we can apply Lemma 5.3 and obtain $\langle(n+1)\alpha\rangle_{*}<\langle n\alpha\rangle_{*}$. Otherwise, if $\langle n_{1}\alpha\rangle_{*}<1/2$, then since $q_{M-1}<m$, the integer $n_{1}$ is in the range [nm+1,n+m1][n-m+1,n+m-1]. By Lemma 5.6, we have $\langle n_{1}\alpha\rangle_{*}<\langle n\alpha\rangle_{*}$, and so $n\neq\widehat{x}_{*}$.
Case 3: L=M1L=M-1 and $n=b_{M-1}q_{M-1}+\sum_{k=M}^{N}b_{k}q_{k}$ with $b_{M-1}\geq 2$. If $L\geq 1$, we set $n_{1}:=n-q_{M-1}>n-m$. Then, by Lemma 5.7, we have $\langle n_{1}\alpha\rangle_{*}<\langle n\alpha\rangle_{*}$, and so $n\neq\widehat{x}_{*}$. If $L=0$, the same follows from Lemma 5.11.
Case 4: $L\leq M-2$. If $L\geq 1$, we set $n_{1}:=n+q_{L+1}<n+m$. Then by Lemma 5.9, we have $\langle n_{1}\alpha\rangle_{*}<\langle n\alpha\rangle_{*}$, and so $n\neq\widehat{x}_{*}$. If $L=0$, the same follows from Lemma 5.10. ∎
The two above lemmas give us a partial result towards the full characterisation in Theorem 2.1; note that the cases (a) and (b) in Lemma 6.5 are almost the same as the cases (i) and (ii) in Theorem 2.1. Next, we provide some lemmas that will be useful in the situation where $\widehat{x}_{*}\neq n$. We start with two lemmas which will, roughly speaking, allow us to assume $\widehat{x}_{*}<n$ without loss of generality.
6.2.  Switching between $n$ and $q_{T}-n$